Due Sunday, Nov 4, 11:59pm EST 9 Calculate pH During Acid Base Titration CONTENT FEEDBACK Question If 37.5 mL of 0.100 M NaOH is added to 10.0 mL of 0.100 M CH,COOH what will be the pH of the resulting solution? CHCOOHaą) + OH (aq) CH,CO (aq) + H,OU) Round your answer to three decimal places. Provide your answer below pi- MORE INSTRUCTION SUBMIT Content attribution
Answer #1

CH3COOH + NaOH ---------> CH3COONa + H2O

millimoles of CH3COOH = 10.0 x 0.10 = 1.0

millimoles of NaOH added = 37.5 x 0.100 = 3.75

3.75 - 1.0 = 2.75 millimoles NaOH left

[NaOH] = 2.75 / 47.5 = 0.056 M

as NaOH is strong base

[NaOH] = [OH-] = 0.056 M

pOH = - log OH-]

pOH = - log [0.056]

pOH = 1.25

pH = 14 - 1.25

pH = 12.75